OpenStudy (anonymous):

i need help plz Solve this non-linear system of equations. Explain your work and show all steps, neatly (please note that the symbol ^ refers to an exponent. The first equation is read as “y equals x squared plus 3x minus 7). Y = x^2 + 3x - 7 3x - y = -2

OpenStudy (campbell_st):

ok... so take the 2nd equation and make y the subject, what do you get...?

OpenStudy (anonymous):

idk

OpenStudy (campbell_st):

ok... so its a case of adding y to both sides of the equation 3x = y + 2 now subtract 2 from both sides of the equation, what do you get..?

OpenStudy (anonymous):

1x = y?

OpenStudy (campbell_st):

nope... oops and my mistake, it should have been add 2 to both sides of the equation 3x + 2 = y so now both equations are in the form y = ..... so you can equate them x^2 + 3x - 7 = 3x + 2 so you have a quadratic equation equal to a linear equation, this means there are possibly 2 points of intersection.. so you now need to solve \[x^2 + 3x - 7 = 3x + 2\] start by addint 7 to both sides of the equation.

OpenStudy (anonymous):

x^2 + 3x = 3x + 9?

OpenStudy (campbell_st):

great now subtract 3x from both sides of the equation

OpenStudy (anonymous):

x^2 = x +9?

OpenStudy (campbell_st):

not quite x^2 + 3x = 3x + 9 - 3x -3x --------------------- x^2 = 9 there are 2 values for x, what are they... take the square root of both sides..

OpenStudy (anonymous):

x^2 = 3

OpenStudy (campbell_st):

not quite \[\sqrt{x^2} = \pm x\] so then using this idea \[x = \pm 3\] or x = 3 and x = -3 to find the ordered pairs substitute each x value into the equation y = 3x + 2 to find the corresponding y values hope it makes sense

OpenStudy (anonymous):

it does